OpenStudy (anonymous):

Find the intercepts and then use them to graph the equation 1.3x - 1.4y= 3.64

OpenStudy (anonymous):

1.3x = 3.64 (x,0) -1.4y = 3.64 (0,y) solve for x and then solve for y

OpenStudy (saifoo.khan):

whn x intercpts y = 0 whn y intercpts x = 0 Now insert & solve.

OpenStudy (anonymous):

I dont know. Honestly guys I was hurt in Afganistan and since I cannot serve in active duty because of all the pain meds I have to take, I am attending college. So please forgive me if it is difficult for me to understand some of the math I am working on because the class they put me in seems far advanced for me

OpenStudy (saifoo.khan):

O, ok.

OpenStudy (saifoo.khan):

x intercpt = 2.8 y intercpt = -2.6